OpenStudy (anonymous):

Simplify x^3+7x^2/x^2+2x-35

OpenStudy (anonymous):

\[x(x^2+2)-28\]

OpenStudy (anonymous):

\[\frac{x^3+7 x^2}{x^2+2 x-35}\text{=}\frac{x^2 (x+7)}{(x-5) (x+7)}\text{=}\frac{x^2}{x-5} \]

OpenStudy (anonymous):

\[x^3+ 7\left(x^2/x^2\right) + 2 x - 35=x^3+2 x-28 \]